OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation x - 2y = 8.

OpenStudy (john_es):

To find x intercept, you must do y=0, and solve for y \[x=8\]To obtain the y intercept, simple x=0, \[-2y=8\Rightarrow y=-4\]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

@John_ES

OpenStudy (john_es):

To find the intercepts you need to find where the line cuts each axis. Do you understand this?

OpenStudy (anonymous):

i think >.<

OpenStudy (john_es):

Then, the "protocol" is the following, - x axis intercept: do y=0, and solve for x. The points that cut the x axis must have y=0. - y axis intercept: do x=0, and solve for y. The points that cut the y axis must have x=0.

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

but my options are x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (anonymous):

i been stuck on my homework can you please help its getting late >.<

OpenStudy (john_es):

Of course, it is the second option the correct one.

OpenStudy (anonymous):

thank you so much for your help do you think you can help on another one ?

OpenStudy (john_es):

Remember, the x intercept must have y=0, so for us is (8,0). And the y intercept must have x=0, so it must be (0,-4).

OpenStudy (john_es):

Yes, tell me.

OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = one fourth; y intercept at (0, 1) slope = one fourth; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (john_es):

First, we rewrite the eqution, \[y=\frac{1}{4}x-1\]And we see it is similar to the general slope form, \[y=mx+n\]where m is the slope and n the y intercept. So, \[m=1/4;\ \ n=-1\]It is the third option.

OpenStudy (anonymous):

thank you very much i honestly wish i could remember how to do the problems i remember while im doing it in class but when i come home im like ?.?

OpenStudy (john_es):

I know it, the best solution is to keep the hard word. ;)